# A NICER View of PSR J1231--1411: A Complex Case

### v1

@@ 5.3 Caveats
The caveats of this work follow mostly those present in the previous NICER analyses as well (see e.g., Bogdanov et al., 2021, for discussion). Additionally, we are now dependent on more restrictive radius priors, e.g., based on the results of Raaijmakers et al. (2021), since the radius could not be constrained fully independently in this work. Therefore, the existence of better solutions far from the R21 prior peak, which is around $12$ km, or at least outside from the [10,14] km range, cannot be ruled out (even though the hard limits for the R21 case were still kept at $\sim 4.4$ and $16$ km).
We also note that our results rely on $M-\cos i$ priors from a preliminary radio timing inference (Cromartie et al., 2024). In the early analyses without these priors, we still found mostly similar (small) masses, but the inclination angles were found to be much smaller, with $\cos i$ peaking between $0.4$ and $0.9$ (depending on the model) instead of at around $0.15$ as predicted by the radio timing. However, these analyses did not yet employ any restricted radius prior for the PDT-U model. We also note that the inclination angle from radio timing refers to the angle between the line of sight and the orbital axis, whereas we use it as the angle between line of sight and NS spin axis. However, previous accretion of matter is expected to have gradually aligned the spin and orbital axes of the system (see, e.g., Bhattacharya & van den Heuvel 1991), and any remaining tilt between them is likely to be insignificant compared to the width of the prior.
We have not tested if setting the lower limit of mass below $1$ $M_{\odot}$ would affect our results. Recent spectral modeling of HESS J1731-347 has suggested that some NSs could have such low masses (Doroshenko et al., 2022). However, this interpretation relies on several assumptions, including the distance associated with the star and the use of a uniform-temperature carbon atmosphere model, as discussed by Alford & Halpern (2023). More importantly, this model is a poor fit to longer, better quality XMM-Newton data from 2014 of the same source, while a two-blackbody model with a mass of, e.g., $1.4$ $M_{\odot}$, provides a good fit to the same data (Alford & Halpern, 2023). Therefore, current data do not support a mass less than $1$ $M_{\odot}$ for HESS J1731-347.
Finally, we note that the inferred temperatures of some of the hot regions are rather cold (notably below $10^{6}$ K), making the effects or partial ionization more relevant in the NS atmosphere, whereas we assume it to be fully ionized both for the hydrogen and helium case. However, the effect is still expected to be minor, since the signal is dominated by the hottest regions. In addition, switching between hydrogen and helium was not found to significantly alter the results (for the R21 case), implying only minor dependence on the atmosphere choice.
@@ 6 Conclusions
We have jointly analyzed NICER and XMM-Newton data for PSR J1231--1411 using X-ray pulse profile modeling. We found that the simple ST-U model with two uniformly emitting spherical caps is not sufficient to explain the data, especially the bolometric pulse profile. Instead, good fits both in bolometric and energy-resolved data and significantly larger Bayesian evidence were found using the PDT-U model, where both hot regions consist of two circular components that are allowed to have different temperatures. Unfortunately, we did not manage to get converged results using a wide NS radius prior with this computationally expensive model. However, forcing the radius to be consistent with previous multi-messenger observations and nuclear theory, we found $R_{\mathrm{eq}}=12.6\pm 0.3$ km and $M=1.04_{-0.03}^{+0.05}$ $M_{\odot}$ and a substantial improvement in evidences and likelihoods. When assuming helium (instead of hydrogen) composition for the NS atmosphere, we were also able reproduce the data well and obtained very similar parameter constraints. Equally good fits, and very similar results were also found for the hydrogen atmosphere when applying an uninformative prior limited between $10$ and $14$ km, except that the radius was inferred to be $R_{\mathrm{eq}}=13.5_{-0.5}^{+0.3}$. In all the best-fitting models we infer a quite similar non-antipodal hot region geometry, implying that a pure dipolar magnetic field is not likely for PSR J1231--1411.
@@ Appendix A Additional Tables and Figures
We summarize the results for the best-fitting hydrogen models (PDT-U H $R_{\mathrm{eq}}\in[10,14]$ km and PDT-U H $R_{\mathrm{eq}}$ R21) in Tables 2 and 3. Additional PDT-U posteriors, including the parameters not shown in Figure 5, are presented in Figures 10, 11, and 12. The main ST-U posterior distributions for radius, compactness and mass are shown in Figure 13.
@@
$\phi_{p}$ [[[[cycles]]]] $p$ superseding component initial phased $\phi_{p}\sim U(-0.25,0.75)$, wrapped $0.38_{-0.01}^{+0.01}$ $5.25$ $0.375$
$\phi_{s}$ [[[[cycles]]]] $s$ superseding component initial phasee $\phi_{s}\sim U(-0.25,0.75)$, wrapped $0.53_{-0.01}^{+0.01}$ $5.29$ $0.536$
@@
non-overlapping hot regions function of all $\Theta$, $\phi$, $\chi$, and $\zeta$
@@
number of processes (multi-modes): $4$
@@
Note. —   We show the prior PDFs, $68.3\,\%$ credible intervals around the median $\widehat{\textrm{CI}}_{68\%}$, KL divergence $\widehat{D}_{\textrm{KL}}$ in bits, and the maximum likelihood nested sample $\widehat{\textrm{ML}}$ for all the parameters. The subscripts $p$ and $s$ denote for primary and secondary hot region parameters, respectively. Note that $\phi_{p}$ ($\phi_{s}$) is measured with respect to the meridian on which Earth (Earth antipode) lies.
a Truncated between $1$ and $3$ $M_{\odot}$. However, note that for each $\cos i$ the $M$ prior is different as seen in Figure 3.
@@
$\phi_{p}$ [[[[cycles]]]] $p$ superseding component initial phased $\phi_{p}\sim U(-0.25,0.75)$, wrapped $0.38_{-0.01}^{+0.01}$ $5.30$ $0.375$
$\phi_{s}$ [[[[cycles]]]] $s$ superseding component initial phasee $\phi_{s}\sim U(-0.25,0.75)$, wrapped $0.53_{-0.01}^{+0.01}$ $5.31$ $0.541$
@@
non-overlapping hot regions function of all $\Theta$, $\phi$, $\chi$, and $\zeta$
@@
number of processes (multi-modes): $4$
@@
Note. —   We show the prior PDFs, $68.3\,\%$ credible intervals around the median $\widehat{\textrm{CI}}_{68\%}$, KL divergence $\widehat{D}_{\textrm{KL}}$ in bits, and the maximum likelihood nested sample $\widehat{\textrm{ML}}$ for all the parameters. The subscripts $p$ and $s$ denote for primary and secondary hot region parameters, respectively. Note that $\phi_{p}$ ($\phi_{s}$) is measured with respect to the meridian on which Earth (Earth antipode) lies.
a Truncated between $1$ and $3$ $M_{\odot}$. However, note that for each $\cos i$ the $M$ prior is different as seen in Figure 3.
@@
c Truncated between $3r_{\rm g}(1)\approx 4.4$ km and $16$ km. The prior density corresponds to the R21 case explained in Section 3.3.2.

### v2

@@ 5.3 Caveats
The caveats of this work mostly follow those present in the previous NICER analyses as well (see e.g., Bogdanov et al., 2021, for discussion). Additionally, we are now dependent on more restrictive radius priors, e.g., based on the results of Raaijmakers et al. (2021), because the radius could not be constrained fully independently in this work. Therefore, the existence of better solutions far from the R21 prior peak, which is around $12$ km, or at least outside from the [10,14] km range, cannot be ruled out (even though the hard limits for the R21 case were still kept at $\sim 4.4$ and $16$ km).
We also note that our results rely on $M-\cos i$ priors from a preliminary radio timing inference (Cromartie et al. 2024, in preparation). In the early analyses without these priors, we still found mostly similar (small) masses but the inclination angles were found to be much smaller, with $\cos i$ peaking between $0.4$ and $0.9$ (depending on the model) instead of at around $0.15$ as predicted by the radio timing. However, these analyses did not yet employ any restricted radius prior for the PDT-U model. We also note that the inclination angle from radio timing refers to the angle between the line of sight and the orbital axis, whereas we use it as the angle between line of sight and NS spin axis. However, previous accretion of matter is expected to have gradually aligned the spin and orbital axes of the system (see, e.g., Bhattacharya & van den Heuvel 1991), and any remaining tilt between them is likely to be insignificant compared to the width of the prior.
We have not tested if setting the lower limit of mass below $1$ $M_{\odot}$ would affect our results. Recent spectral modeling of HESS J1731--347 has suggested that some NSs could have such low masses (Doroshenko et al., 2022). However, this interpretation relies on several assumptions, including the distance associated with the star and the use of a uniform-temperature carbon atmosphere model, as discussed by Alford & Halpern (2023). More importantly, this model is a poor fit to longer, better quality XMM-Newton data from 2014 of the same source, while a two-blackbody model with a mass of, e.g., $1.4$ $M_{\odot}$, provides a good fit to the same data (Alford & Halpern, 2023). Therefore, current data do not support a mass less than $1$ $M_{\odot}$ for HESS J1731--347.
Finally, we note that the inferred temperatures of some of the hot regions are rather cold (notably below $10^{6}$ K), making the effects or partial ionization more relevant in the NS atmosphere, whereas we assume it to be fully ionized both for the hydrogen and helium case. However, the effect is still expected to be minor because the signal is dominated by the hottest regions. In addition, switching between hydrogen and helium was not found to significantly alter the results (for the R21 case), implying only minor dependence on the atmosphere choice.
@@ 6 Conclusions
We have jointly analyzed NICER and XMM-Newton data for PSR J1231--1411 using X-ray pulse profile modeling. We found that the simple ST-U model with two uniformly emitting spherical caps is not sufficient to explain the data, especially the bolometric pulse profile. Instead, good fits both in bolometric and energy-resolved data and significantly larger Bayesian evidence were found using the PDT-U model, where both hot regions consist of two circular components that are allowed to have different temperatures. Unfortunately, we did not manage to get converged results using a wide NS radius prior with this computationally expensive model. However, forcing the radius to be consistent with previous multimessenger observations and nuclear theory, we found $R_{\mathrm{eq}}=12.6\pm 0.3$ km and $M=1.04_{-0.03}^{+0.05}$ $M_{\odot}$ and a substantial improvement in evidence and likelihoods. When assuming helium (instead of hydrogen) composition for the NS atmosphere, we were also able reproduce the data well and obtained very similar parameter constraints. Equally good fits, and very similar results, were also found for the hydrogen atmosphere when applying an uninformative prior limited between $10$ and $14$ km, except that the radius was inferred to be $R_{\mathrm{eq}}=13.5_{-0.5}^{+0.3}$ km. In all the best-fitting models we infer a quite similar nonantipodal hot region geometry, implying that a pure dipolar magnetic field is not likely for PSR J1231--1411.
Acknowledgments
@@ Appendix A Additional Tables and Figures
We summarize the results for the best-fitting hydrogen models (PDT-U H $R_{\mathrm{eq}}\in[10,14]$ km and PDT-U H $R_{\mathrm{eq}}$ R21) in Tables 2 and 3. Additional PDT-U posteriors, including the parameters not shown in Figure 5, are presented in Figures 10, 11, and 12. The main ST-U posterior distributions for radius, compactness, and mass are shown in Figure 13.
@@
$\phi_{p}$ [[[[cycles]]]] $p$ superseding component initial phase $\phi_{p}\sim U(-0.25,0.75)$, wrapped $0.38_{-0.01}^{+0.01}$ $5.25$ $0.375$
$\phi_{s}$ [[[[cycles]]]] $s$ superseding component initial phase $\phi_{s}\sim U(-0.25,0.75)$, wrapped $0.53_{-0.01}^{+0.01}$ $5.29$ $0.536$
@@
Nonoverlapping hot regions function of all $\Theta$, $\phi$, $\chi$, and $\zeta$
@@
Number of processes (multimodes): $4$
@@
Note. —   We show the prior PDFs, $68.3\,\%$ credible intervals around the median $\widehat{\textrm{CI}}_{68\%}$, KL-divergence $\widehat{D}_{\textrm{KL}}$ in bits, and the maximum-likelihood nested sample $\widehat{\textrm{ML}}$ for all the parameters. The subscripts $p$ and $s$ denote for primary and secondary hot region parameters, respectively. Note that $\phi_{p}$ ($\phi_{s}$) is measured with respect to the meridian on which Earth (Earth antipode) lies.
a Truncated between $1$ and $3$ $M_{\odot}$. However, note that for each $\cos i$ the $M$ prior is different, as seen in Figure 3.
@@
$\phi_{p}$ [[[[cycles]]]] $p$ superseding component initial phase $\phi_{p}\sim U(-0.25,0.75)$, wrapped $0.38_{-0.01}^{+0.01}$ $5.30$ $0.375$
$\phi_{s}$ [[[[cycles]]]] $s$ superseding component initial phase $\phi_{s}\sim U(-0.25,0.75)$, wrapped $0.53_{-0.01}^{+0.01}$ $5.31$ $0.541$
@@
Nonoverlapping hot regions function of all $\Theta$, $\phi$, $\chi$, and $\zeta$
@@
Number of processes (multimodes): $4$
@@
Note. —   We show the prior PDFs, $68.3\,\%$ credible intervals around the median $\widehat{\textrm{CI}}_{68\%}$, KL-divergence $\widehat{D}_{\textrm{KL}}$ in bits, and the maximum-likelihood nested sample $\widehat{\textrm{ML}}$ for all the parameters. The subscripts $p$ and $s$ denote for primary and secondary hot region parameters, respectively. Note that $\phi_{p}$ ($\phi_{s}$) is measured with respect to the meridian on which Earth (Earth antipode) lies.
a Truncated between $1$ and $3$ $M_{\odot}$. However, note that for each $\cos i$ the $M$ prior is different, as seen in Figure 3.
@@
c Truncated between $3r_{\rm g}(1)\approx 4.4$ km and $16$ km. The prior density corresponds to the R21 case that was explained in Section 3.3.2.